# ImageSet2Text: Describing Sets of Images through Text

### v1

@@ C Completeness Evaluation
To assess the completeness of our descriptions, we conduct an experiment on the downstream task of Set Difference Captioning using the PairedImageSets dataset [18].
@@
First, the input provided to the proposer to identify differences between the sets is the graph representations generated at the final iteration of ImageSet2Text for both image sets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. These graphs are then transformed into a textual format using the generate_network_text function from the NetworkX library in Python [^6] NetworkX, https://networkx.org/documentation/stable/reference/readwrite/generated/networkx.readwrite.text.generate_network_text.html, Last Access: 7th of March 2025..
In the original VisDiff implementation, the authors conduct three rounds of their pipeline, where in each round, 10 different images are considered to generate 10 candidate differences. The proposed differences are merged over the three rounds (for a total of 30 differences) and then passed as input to the ranker module. Since our graph representations are precomputed and remain static, we have considered using a single round of iteration. However, we observed that when the proposer is prompted to find 30 possible differences at once, the proposals start becoming meaningless, diverging toward irrelevant interpretations. To mitigate this, we adopt a two-round approach, requesting 15 differences per round, for a total of 30 proposed differences.
@@
In the main paper, we demonstrated that integrating the information extracted through ImageSet2Text enhances performance on the PairedImageSets dataset. In this section, we further examine this improvement through a case-by-case analysis, directly comparing our results with those of VisDiff on six specific failure cases reported in their paper [18]. The comparison is illustrated in Fig. 8.
@@
D User Study
To assess the quality of descriptions generated by ImageSet2Text, we conducted a user study with 200 participants recruited via Prolific. The study was implemented as a dynamic Google Form, deployed as a Google Web Application, and coded using Google Apps Script. Each user evaluates seven descriptions: six generated by ImageSet2Text and one control description. Three examples of control descriptions (one per type) and their corresponding image grids are shown in Fig. 9. The study takes approximately 8 minutes per user. The user study questions are listed in Tab. 8 and a screenshot of the interface is reported in Fig. 10, where a pair set-description is reported, along with the first question of the user study.
@@ D User Study
E Automatic Alternative Text Generation of Image Sets
@@
Automatic generation of alternative text (alt-text) is a fundamental application of image captioning [22], particularly for visually impaired individuals. ImageSet2Text introduces a novel approach by summarizing entire image collections rather than generating captions for individual images. Since this area is largely unexplored, we conducted interviews with three visually impaired individuals by virtue of a collaboration with Fundación ONCE, a Spanish NGO devoted to improving universal accessibility. The goal of the interviews was to gather community feedback on its usefulness and potential improvements [13]. The interviews were conducted by the first author in Spanish. They were transcribed and then carefully translated into English to facilitate a qualitative analysis of the collected data. The questions are summarized in Tab. 9.
@@
Interviewees generally welcomed the idea, recognizing that set-level descriptions could be extremely useful when understanding the broader context of a scene is more important than focusing on specific details, such as during events and entertainment, keeping memories of travels, or while managing folders on the computer. However, they emphasized that such summaries should complement rather than replace individual image descriptions, as both serve distinct purposes. When evaluating an example description, participants expressed general satisfaction with the level of detail, coherence, and clarity. In particular, they appreciated explicit relations among the entities in the images—an aspect they often find lacking in commercial automatic alt-text generators. This feature of ImageSet2Text is a direct consequence of integrating structural representations that explicitly consider relationships between visual elements [49, 50].
Should ImageSet2Text be further developed in the context of accessible technologies, our collaborators from Fundación ONCE suggested key areas for improvement. First, they emphasized the importance of using simple, clear, and direct language to minimize ambiguity. They also recommended tailoring descriptions based on the user’s visual experience—for instance, those who have seen before might benefit from references to colors and light, while those who have never seen may require alternative descriptions. Another point raised was that the current approach works best for homogeneous image sets with shared visual elements. However, in real-world scenarios, image collections might be more heterogeneous. As a result, a necessary future direction for ImageSet2Text is to not only identify common features but also to detect distinct groups within an image set to provide more meaningful summaries, which is aligned with ongoing research in semantic image clustering [38].
In sum, this feedback highlights the potential of ImageSet2Text to enhance accessibility and inclusion for blind and low-vision users in both personal and professional settings.

### v2

@@ Appendix C Completeness Evaluation
To assess the completeness of our descriptions, we conduct an experiment on the downstream task of Set Difference Captioning using the PairedImageSets dataset (visdiff).
C.1 Integrating ImageSet2Text into VisDiff
@@
First, the input provided to the proposer to identify differences between the sets is the concept graph representations generated at the final iteration of ImageSet2Text for both image sets $\mathcal{D}_{A}$ and $\mathcal{D}_{B}$. The concept graphs are transformed into a textual format using the generate_network_text function from the NetworkX library in Python [^9] NetworkX, https://networkx.org/documentation/stable/reference/readwrite/generated/networkx.readwrite.text.generate˙network˙text.html, Last Access: 7th of March 2025..
In the original VisDiff implementation, the authors conduct three rounds of their pipeline, where, in each round, 10 different images are considered to generate 10 candidate differences. The proposed differences are merged over the three rounds (for a total of 30 differences) and then passed as input to the ranker module. Since our concept graphs are precomputed and remain static, we have considered using a single round of iteration. However, we observed that when the proposer is prompted to find 30 possible differences at once, the proposals start becoming meaningless, diverging toward irrelevant interpretations. To mitigate this, we adopt a two-round approach, requesting 15 differences per round, for a total of 30 proposed differences.
@@
C.2 Failure cases of VisDiff
In the main paper, we demonstrated that integrating the information extracted through ImageSet2Text enhances performance on the PairedImageSets dataset. In this section, we further examine this improvement through a case-by-case analysis, directly comparing our results with those of VisDiff on six specific failure cases reported in their paper (visdiff). The comparison is illustrated in Fig. 12.
@@ C.2 Failure cases of VisDiff
C.3 Compute Infrastructure
The experiments investigated in this section were conducted in Google Colab using a Tesla T4 GPU (16GB VRAM). Calls to OpenAI GPT models were performed using the latest API version as of March 2025. We provide the full code of our experiments, with the most important packages being:
  • python = 3.11
  • pytorch = 2.6.0+cu124
  • nltk = 3.9.1
Appendix D User Study
To assess the quality of descriptions generated by ImageSet2Text, we conducted a user study with 233 participants recruited via Prolific. The study was implemented as a dynamic Google Form, deployed as a Google Web Application, and coded using Google Apps Script. Each user evaluates seven descriptions: six generated by ImageSet2Text and one control description. Three examples of control descriptions (one per type) and their corresponding image grids are shown in Fig. 13. The study takes approximately 8 minutes per user. The user study questions are listed in Table 8 and a screenshot of the interface is reported in Fig. 14, where a pair set-description is reported, along with the first question of the user study.
@@ Appendix D User Study
Appendix E Failure cases
In this section, we detail the failure cases caused by the integration of WordNet and CLIP in the verification process of ImageSet2Text, along with an experimental analysis to estimate their frequency and impact on the quality of the generated descriptions. We also discuss additional limitations inherited from the individual components that constitute the method’s pipeline.
(a) Four examples of nodes with siblings that are not mutually exclusive. From left to right, the first two fall into the rhomboid case and the second two fall into the triangular case.
(b) Four examples of sibling nodes that are semantically distinct but visually too similar to be effectively distinguished in the CLIP embedding space.
Figure 15: Illustration of failure cases involving non-mutually-exclusive or visually similar sibling nodes in WordNet.
E.1 Non-mutually exclusive sibling nodes
As described in Section 3 of the main paper, given a hypothesis $h_{i}\in H$ to verify, the set of contradicting hypotheses $\mathcal{H}_{i}^{-}$ is constructed by replacing the object $o_{i}$ in $h_{i}$ with its sibling nodes. This approach assumes that sibling nodes are mutually exclusive: a condition that does not always hold in WordNet. Some representative counterexamples are shown in Fig. 15(a), which illustrates two distinct structural patterns:
  • Rhomboid case: A node has two hyponyms that share a common descendant. For example, proboscidean and pachyderm are sibling nodes, but both have elephant as a hyponym, and are therefore not mutually exclusive. A similar structure is seen with tennis player, which is a child of both participant and athlete.
  • Triangular case: A lexical concept can be linked to a hypernym at different levels of granularity. For instance, limestone is a child of both rock and sedimentary rock, where the latter is itself a child of rock. Consequently, while limestone and sedimentary rock appear as siblings under rock, they are also hierarchically connected and not mutually exclusive. A similar case is observed with hand shovel, as shown in the figure.
The existence of these structural patterns poses challenges for the verification phase. Therefore, it is important to estimate how many nodes fall under this condition to its impact in the definition of the set $\mathcal{H}_{i}^{-}$. To estimate the number of affected nodes, we define as intersection leaf nodes those that have multiple hypernyms in the WordNet hierarchy. Given an intersection node, we trace upwards through the hierarchy to identify their lowest common ancestor, which we refer to as the intersection root node (e.g., eutherial mammal is the intersection root node for the intersection leaf node elephant).
The verification issue arises for the hyponyms of an intersection root node, since at least two of them lead to the intersection leaf node, hence representing concepts that are not mutually exclusive. We compute the total number of intersection root nodes in WordNet and compare that to the total number of nodes eligible for verification (i.e., all non-leaf nodes). Our analysis shows that only 1.49% of the nodes are affected by this issue. This low percentage supports the use of the current definition of $\mathcal{H}_{i}^{-}$ in our extensive experiments, where it has not posed significant problems. However, despite its rarity, this failure case should be acknowledged in future iterations of this work, where more robust strategies could be considered.
E.2 Visual similarity among siblings
The second failure case that we identified also affects the verification phase. Specifically, it involves the kNN classifier operating over sibling nodes that represent mutually exclusive concepts which are not visually distinguishable. Representative examples are shown in Fig. 15(b). In such cases, the kNN classifier may struggle to differentiate between the hypothesis $h_{i}$ under verification and its contradicting hypotheses in $\mathcal{H}_{i}^{-}$. While this limitation is acceptable in general (i.e., the system only needs to be as sensitive as the visual data allows), it becomes an issue when these indistinguishable siblings are embedded within a broader class that includes clearly distinct concepts.
For instance, suppose we are verifying the hypothesis that a set of images depicts a tea-like drink, but the hypothesis is rejected due to the presence of tea among the contradicting alternatives, which is visually inseparable from tea-like drink in the CLIP embedding space. As a result, the verification process halts at a higher-level concept in the lexical hierarchy, such as beverage, which is too generic to capture the intended meaning. This loss of granularity leads to the generation of overly vague descriptions, even when the visual evidence could support more precise interpretations.
To assess the severity of this issue, we perform an empirical estimate of how frequently this failure case occurs and how significantly it impacts the verification process. Unlike the previous case, this analysis requires access to visual information, which is not provided by WordNet. Interestingly, ImageNet (deng2009) builds its classes based on the WordNet nouns hierarchy, making it a suitable resource for investigating visually indistinguishable sibling nodes. It is important to note that not all WordNet nodes are present in ImageNet, as some concepts are inherently non-visual and therefore excluded. For our analysis, we rely on a subset of ImageNet, assuming that the resulting percentage of affected cases is representative of what would be observed over the full dataset. In particular, over the total of 16693 nodes eligible for verification (i.e. the non-leaf nodes), only 5991 has at least one hyponym present in ImageNet. Because we had access to a reduced version of ImageNet, we consider 874 of these nodes.
For each node (e.g. beverage) we consider all possible hyponym pairs, which can be either visually distinguishable (e.g. tea and coffee) or visually indistinguishable (e.g. tea and tea-like drink). We define a pair visually distinguishable if, given five images per hyponym, CLIP correctly matches at least 8 out of 10 images to their respective hyponym, based on similarity scores. Conversely, we define the pair visually indistinguishable if this condition does not hold. In addition, we consider a node as overgeneric if the proportion of visually indistinguishable pairs of its hyponyms falls below a predefined threshold. Formally, a node is marked as overgeneric if:
$\frac{\#_{\text{indist}}}{\#_{\text{dist}}+\#_{\text{indist}}}<\phi$
where:
  • $_{\text{dist}}$ is the number of visually distinguishable hyponym pairs of the node,
  • $_{\text{indist}}$ is the number of visually indistinguishable hyponym pairs of the node,
  • and $\phi$ is a predefined threshold.
Using this criterion, and setting $\phi=0.15$ we identified 51 unreliable nodes, corresponding (in proportion) to $2.09\%$ of the total non-leaf nodes of WordNet, providing an estimate of the prevalence of this specific failure case. By setting the threshold $\phi$ to a low value, we capture situations where failures emerge not from CLIP’s inherent weaknesses (discussed in Section E.3), but from the presence of concepts actually indistinguishable visually.
Given the low occurrence of this failure case, its impact is negligible in the current version of ImageSet2Text. However, in future iterations of this work, it may be beneficial to account for such ambiguity. One possible solution could involve introducing disjunction nodes, i.e. allowing ImageSet2Text to express uncertainty through formulations like tea or tea-like drink.
E.3 Other Limitations
In addition to the previously discussed failure cases, primarily arising from the integration of WordNet and CLIP during the verification phase, we identify further limitations of ImageSet2Text that stem from the individual components of the pipeline.
One notable constraint concerns the language model: ImageSet2Text currently depends on GPT-4o-mini, a proprietary model. In future work, we aim to investigate open-source alternatives to enhance the accessibility of the pipeline for researchers and users with limited resources.
As a pipeline that heavily relies on large language models, ImageSet2Text also inherits many of their well-documented limitations. These include a tendency to generate plausible but incorrect content, sensitivity to prompt phrasing, and difficulties with factual consistency and reasoning in complex tasks (bommasani2021opportunities). Broader challenges, such as opacity, and the risk of biased or unreliable outputs, are especially salient in foundation models (liang2023holistic). Our prompt engineering efforts were devoted to mitigate these limitations as much as possible. As research on LLMs progresses and these limitations are addressed by the community, also the quality of the descriptions generated by ImageSet2Text is expected to improve.
We also highlight limitations related to the usage of CLIP as a zero-shot classifier for verification. While CLIP has proven remarkably effective for this task over broad and generic categories, its performance degrades significantly in fine-grained and open-world settings (bianchi2024clip). A central limitation lies in CLIP’s difficulty distinguishing between semantically or visually similar classes, especially when these distinctions are subtle or domain-specific. This stems from both the model’s reliance on internet-scale, weakly labeled data and its coarse-grained alignment between images and text. These limitations suggest that CLIP, while powerful, remains a bottleneck for open-world applications of ImageSet2Text.
Another limitation we encountered arises during the hypothesis expansion phase, where reliance on WordNet introduces ambiguity due to polysemy: i.e., the presence of multiple definitions for the same word, as is common in any lexical resource. In some cases, this can lead to mismatches between the intended definition of a word and the WordNet selected hypernyms for generalizing the hypothesis $o_{0}$ in $h_{0}$ and generating the set $\mathcal{H}$. We were able to mitigate this issue effectively by prompting the LLM to select the most contextually appropriate WordNet node, based on $o_{0}$. This simple yet effective strategy allowed us to maintain semantic precision in the expansion process without requiring manual intervention.
Appendix F Automatic Alternative Text Generation of Image Sets
@@
Automatic generation of alternative text (alt-text) is a fundamental application of image captioning (gurari2020), particularly for visually impaired individuals. ImageSet2Text introduces a novel approach by summarizing entire image collections rather than generating captions for individual images. Since this area is largely unexplored, we conducted interviews with three collaborators from the Spanish national association for universal accessibility, Fundación ONCE, to gather community feedback on its usefulness and potential improvements (costanza2020). The interviews were conducted by the first author in Spanish. They were transcribed and then carefully translated into English to facilitate a qualitative analysis of the collected data. The questions are summarized in Table 9.
@@
Our collaborators generally welcomed the idea, recognizing that set-level descriptions could be extremely useful when understanding the broader context of a scene is more important than focusing on specific details, such as during events and entertainment, keeping memories of travels, or while managing folders on the computer. However, they emphasized that such summaries should complement rather than replace individual image descriptions, as both serve distinct purposes. When evaluating an example description, participants expressed overall satisfaction with the level of detail, coherence, and clarity. In particular, they appreciated explicit relations among the entities in the images, an aspect they often find lacking in commercial automatic alt-text generators. This feature of ImageSet2Text is a direct consequence of integrating structural representations that explicitly consider relationships between visual elements (phueaksri2023; phueaksri2024).
Should ImageSet2Text be further developed in the context of accessible technologies, our collaborators from Fundación ONCE suggested key areas for improvement. First, they emphasized the importance of using simple, clear, and direct language to minimize ambiguity. They also recommended tailoring descriptions based on the user’s visual experience, e.g. those who have seen before might benefit from references to colors and light, while those who have never seen may require alternative descriptions. Another point raised was that the current approach works best for homogeneous image sets with shared visual elements. In real-world scenarios, however, image collections might be more heterogeneous. As a result, a necessary future direction for ImageSet2Text is to not only identify common features but also detect distinct clusters within an image set to provide more meaningful summaries, aligning with ongoing research in semantic image clustering (liu2024).
Overall, this feedback highlights the potential of ImageSet2Text to enhance accessibility and inclusion for blind and low-vision users in both personal and professional settings.